# Second Law of Entanglement Manipulation with Entanglement Battery

### v1

@@ Abstract
A central question since the beginning of quantum information science is how two distant parties can convert one entangled state into another. Answers to these questions enable us to optimize the performance of tasks such as quantum key distribution and quantum teleportation, since certain entangled states are more useful than others for these applications. It has been conjectured that entangled state transformations could be executed reversibly in an asymptotic regime, mirroring the reversible nature of Carnot cycles in classical thermodynamics. While a conclusive proof of this conjecture has been missing so far, earlier studies excluded reversible entanglement manipulation in various settings. In this work, we investigate the concept of an entanglement battery, an auxiliary quantum system that facilitates quantum state transformations without a net loss of entanglement. We establish that reversible manipulation of entangled states is achievable through local operations when augmented with an entanglement battery. In this setting, two distant parties can convert any entangled state into another of equivalent entanglement. The rate of asymptotic transformation is quantitatively expressed as a ratio of the entanglement present within the quantum states involved. Different entanglement quantifiers give rise to unique principles governing state transformations, effectively constituting diverse manifestations of a “second law” of entanglement manipulation. Our methods provide a solution to the long-standing open question regarding the reversible manipulation of entangled states and are also applicable to entangled systems involving more than two parties, and to other quantum resource theories, including quantum thermodynamics.
The industrial revolution was a transformative era spanning the 18th and 19th centuries. It was significantly driven by breakthrough discoveries in statistical physics and thermodynamics. These disciplines provided crucial insights into energy conversion, particularly in heat engines that paved the way for advancements in technology and industry. Today, we stand on the cusp of a similar threshold, this time driven by quantum technology. This field hinges on the intricate properties of quantum systems, with quantum entanglement [1] and coherence [2] playing a pivotal role.
Over the past decades, striking parallels between the principles governing manipulation of entangled systems and the laws of thermodynamics [3, 4, 5, 6, 7] have been revealed. A prime illustration of this similarity is often cast with the narrative of two agents, Alice and Bob, sharing $n$ copies of an entangled state |$\ket{\psi}$ that they can also manipulate. It is known that under certain conditions, by utilizing simple operations such as local operations and classical communication [8], Alice and Bob can transform their initially shared state into $n$ copies of another state |$\ket{\phi}$. In the regime of large $n$, this transformation is contingent upon an important condition: a reduction in entanglement entropy [8]. This rule mirrors a fundamental concept in classical thermodynamics, where the entropy of a system uniquely determines its potential for interconversion through adiabatic processes [9]. This resemblance highlights the universality of entropy as a key concept in both domains, for understanding and describing state transformations.
The striking similarity between entanglement theory and thermodynamics naturally leads to an intriguing inquiry: Does there exist a “second law of entanglement manipulation” [3], akin to its thermodynamic counterpart, that governs all state transformations of entangled systems? This question is linked to the feasibility of reversibility in quantum mechanics, in particular when it comes to manipulating entangled states in an asymptotic setting. This enables a theoretically lossless conversion between any two entangled states in the asymptotic limit. Furthermore, this mirrors classical thermodynamics, where Carnot’s theorem connects the reversibility of a heat engine cycle and its efficiency.
Despite numerous dedicated efforts [3, 4, 5, 6, 7, 10, 11], the quest for a second law of entanglement manipulation is still ongoing. Ref. [7] establishes that there is no second law of entanglement manipulation through a certain class of deterministic protocols. However, this no-go statement can still be bypassed via relaxations such as probabilistic protocols [11]. One of the main challenges in this endeavor is the existence of bound entanglement [12]. Bound entangled quantum states, akin to black holes in the realm of astrophysics, require entanglement for their formation, yet they defy the extraction of usable entanglement in the form of singlets [12]. This paradoxical feature of entanglement poses significant theoretical and practical challenges, complicating the path towards a fully reversible framework of entangled state manipulations.
@@
In this article, we resolve this conundrum by focusing on state transformations instead of protocols. We introduce the concept of entanglement batteries [13] into our framework. An entanglement battery is an additional entangled quantum system shared between Alice and Bob. In order to prevent the embezzling of resources [14], the battery must be returned at the end of the procedure with at least the same amount of entanglement. This idea can be viewed as a generalization of catalytic state transformations [15], a topic that has garnered significant interest recently [16, 17]. An entangled catalyst, in this context, is an ancillary quantum system in an entangled state provided to Alice and Bob. They are allowed to employ this catalyst in their transformation process, with the requirement that it must be returned in its original state. Our approach extends this setup by allowing changes in the state of the ancillary system, provided that there is no reduction in its entanglement.
We demonstrate in this work that such an approach leads to a second law of entangled state manipulation. This law asserts that Alice and Bob can transform a state $\rho$ into another $\sigma$ if and only if $E(\rho)\geq E(\sigma)$, where $E$ represents an entanglement quantifier with certain natural properties. In the asymptotic limit, the conversion rates in this setting are characterized by the ratio $E(\rho)/E(\sigma)$, leading to the implication that asymptotic state manipulations can be executed in a reversible manner. Our results lead to a family of second laws of entanglement manipulation, contingent on the method of entanglement quantification in both the system and the battery. Moreover, we recover a variant of the second law previously conjectured in [4], achieved without resorting to the generalized quantum Stein’s lemma, for which a comprehensive proof is still missing [10].
Asymptotic transformations of entangled states and reversibility
@@
A fundamental challenge in quantum information science revolves around the interconversion of quantum states by two distant agents, Alice and Bob [18, 1]. Given a shared entangled quantum state, the question arises: What quantum states can Alice and Bob obtain, if they are restricted to local operations and classical communication (LOCC)? By LOCC, we mean local transformations within their respective laboratories and communication via classical messages. Addressing this problem is crucial for delineating the utility of specific quantum states within this framework, effectively identifying the states that offer the greatest utility for quantum information processing and communication tasks.
The problem posed above is closely connected to the important process of entanglement distillation, which involves extracting singlets |$\ket{\psi^{-}}=(\ket{01}-\ket{10})/\sqrt{2}$ from a – possibly noisy – quantum state [19]. Singlets are “gold standards” of entanglement, which have a pivotal role in foundational tasks such as quantum teleportation [20] and quantum key distribution [21]. Typically, entanglement distillation is explored within an asymptotic framework, wherein Alice and Bob have access to $n$ copies of a quantum state $\rho$, and their aim is to generate $m$ singlets. A key metric of success in this endeavor is the maximum achievable ratio $m/n$. An error margin is permitted in the transformation, with the stipulation that the error vanishes in the limit of large $n$.
Provided with a large number of copies of a pure entangled state |$\ket{\psi}$, Alice and Bob have the capability to transform these into singlets at a rate determined by the entanglement entropy $E(\ket{\psi})=S(\psi^{A})$ [8], where $S(\rho)=-\mathrm{Tr}(\rho\log_{2}\rho)$ is the von Neumann entropy. Expanding upon this, Alice and Bob are also able to convert |$\ket{\psi}$ into another entangled state |$\ket{\phi}$ at a rate of $S(\psi^{A})/S(\phi^{A})$ [8]. Consequently, this framework enables a reversible, asymptotic conversion between |$\ket{\psi}$ and |$\ket{\phi}$, allowing for a lossless interconversion in the limit of large numbers of copies.
In general, a transformation between two entangled states $\rho$ and $\sigma$ is deemed reversible if the asymptotic transformation rates satisfy the relation $R(\rho\rightarrow\sigma)\times R(\sigma\rightarrow\rho)=1$. However, LOCC manipulations of entangled states generally do not exhibit reversibility, meaning that there exist states with the property $R(\rho\rightarrow\sigma)\times R(\sigma\rightarrow\rho)<1$. An extreme manifestation of such irreversibility arises from the phenomenon of bound entanglement [12]. This phenomenon is characterized by entangled states $\rho$ from which singlets cannot be extracted, such that $R(\rho\rightarrow\psi^{-})=0$, while $R(\psi^{-}\rightarrow\rho)>0$, indicating that singlets are still necessary for the state’s formation [22]. This disparity signifies that the conversion between $\rho$ and |$\ket{\psi^{-}}$ cannot be performed in a reversible manner.
The findings presented in Ref. [4] indicate that reversibility of entangled state transformations may be achievable by expanding the set of LOCC to encompass protocols that permit the injection of a small amount of entanglement into the system, which vanishes in the asymptotic limit. Nonetheless, it is important to note that the assertions in [4] rely on the generalized quantum Stein’s lemma [5]. Recent scrutiny, as discussed in [10], has raised questions about the validity of this lemma, casting uncertainty on the robustness of the conclusions drawn from its application. On the other hand, reversibility within this framework can be achieved probabilistically [11].
Recent studies have also established that reversibility cannot be achieved in settings limited to non-entangling operations, which are protocols incapable of generating entanglement from non-entangled states [7]. Furthermore, reversible manipulations of entangled states cannot be obtained by using entanglement catalysis due to the existence of bound entanglement in this setting [23, 24].
Reversible entanglement manipulations with entanglement battery
@@
We demonstrate that reversible manipulations across all entangled states become feasible when an entanglement battery is incorporated. Specifically, we consider a setting where Alice and Bob can perform LOCC, and additionally are given access to a supplementary shared entangled state, which is our entanglement battery. This setting has been previously studied in Ref. [13], where it was shown that we can formulate a fluctuation theorem in pure state entanglement transformations with an explicit battery model. It was shown that the average change in the battery entanglement when transforming a pure state $\psi^{AB}$ to $\phi^{AB}$ is bounded above by the change in entanglement $S(\psi_{A})-S(\phi_{A})$. In this article, we generalize this framework to allow for a general battery and we study transformations between general mixed states. We show that we obtain a reversible theory when we impose the following restriction on the battery: Alice and Bob must not decrease the level of entanglement within the battery throughout the process. The setup is shown in Fig. 1.
@@
The basic property of any entanglement measure $E$ is that it should not increase under LOCC, i.e., $E(\Lambda[\rho])\leq E(\rho)$ for any LOCC protocol $\Lambda$ [25]. A particularly interesting class of entanglement measures for our analysis are quantifiers which are additive, i.e., they satisfy $E(\rho\otimes\tau)=E(\rho)+E(\tau)$ for any states $\rho$ and $\tau$. Additionally, we will consider entanglement quantifiers that exhibit asymptotic continuity [26]. This property implies that for two states close to each other in trace distance, the difference in the amount of entanglement can grow at most logarithmically in the Hilbert space dimension. We further require the entanglement measure to be finite for all states.
The subsequent theorem offers a comprehensive characterization of state transformations within this setup.
@@
We provide a concise outline of our proof, with comprehensive details available in the Methods section. First, we leverage the properties of additive entanglement measures to establish that under the framework of LOCC augmented with an entanglement battery, the amount of entanglement does not increase. Subsequently, we introduce a protocol that transforms the state $\rho$ into $\sigma$ for any pair of states that comply with the stipulations outlined in Eq. (3). This protocol is remarkably straightforward: the battery is initialized in the target state $\sigma$, and then local transformations are employed to interchange the states of the main system and the battery. A similar technique has been used recently in [17] in the context of correlated catalysis, see also [13].
@@
An outline of the proof is provided here, with detailed elaboration available in the Methods section. Employing the properties of additive and asymptotically continuous entanglement measures, we first demonstrate that the transformation rate is upper-bounded by $E(\rho)/E(\sigma)$. The converse is shown by introducing a protocol that accomplishes the transformation at the aforementioned rate. Echoing the approach of Theorem 1, the optimal protocol involves a battery in an entangled state, specifically comprising copies of the target state $\sigma$. The transformation $\rho\rightarrow\sigma$ at the rate of $E(\rho)/E(\sigma)$ can be realized by permuting the main system with the battery.
An immediate consequence of our work is the construction of a framework of reversible entanglement manipulation. More specifically, we establish that $R(\rho\rightarrow\sigma)\times R(\sigma\rightarrow\rho)=1$ for any $\rho$ and $\sigma$, implying that in the asymptotic limit, any two entangled states $\rho$ and $\sigma$ can be interconverted reversibly. Furthermore, when the ratio $E(\rho)/E(\sigma)$ is rational, reversible interconversion is feasible even with a finite number of copies. This means there exist integers $m$ and $n$ such that $m/n=E(\rho)/E(\sigma)$, allowing for bidirectional conversion $\rho^{\otimes n}\rightarrow\sigma^{\otimes m}$ and $\sigma^{\otimes m}\rightarrow\rho^{\otimes n}$. We further notice that the optimal conversion rate can be achieved through a protocol involving local operations alone, which interestingly obviates the need for classical communication in this task. We suspect that this arises from allowing the battery to contain any finite amount of entanglement. However, we anticipate that achieving the conversion with minimal entanglement in the battery may require some level of classical communication.
@@
with the quantum conditional mutual information $I(A;B|E)$. We refer to the Methods section for more details about the properties of squashed entanglement. We note that the squashed entanglement is not the only entanglement quantifier having these properties, another example is given in the Methods section.
Our findings not only affirm the existence of a second law for entangled state transformations, but also link it to the squashed entanglement present in the involved quantum states. Prior to our work, it was hypothesized [4] that should reversible manipulation of entangled states be achievable in any framework, the rates at which such reversible transformations occur would be linked to a different measure of entanglement: the regularized relative entropy of entanglement. This quantity is defined as [25]
@@
As we will see in the following, our methodologies extend to establish a version of the second law of entanglement manipulations predicated on the relative entropy of entanglement. This can be achieved by replacing the squashed entanglement by $E_{\mathrm{r}}^{\infty}$ in our approach, i.e., the requirement that the regularized relative entropy of entanglement in the battery does not decrease during the procedure. It is currently unknown whether $E_{\mathrm{r}}^{\infty}$ is additive for all states, which prevents a direct application of our results. If $E_{\mathrm{r}}^{\infty}$ exhibits additivity, our results directly imply that $E_{\mathrm{r}}^{\infty}(\rho)/E_{\mathrm{r}}^{\infty}(\sigma)$ is the optimal transformation rate in this setting. Nevertheless, in the Methods section we show that this configuration results in a meaningful entanglement theory with finite asymptotic distillation rates, even if $E_{\mathrm{r}}^{\infty}$ is not additive in general. Moreover, we show that $E_{\mathrm{r}}^{\infty}(\rho)/E_{\mathrm{r}}^{\infty}(\sigma)$ is an achievable rate for the transformation $\rho\rightarrow\sigma$ in this setting. Intriguingly, this closely resembles the second law of entanglement manipulation as delineated in [4], yet it circumvents the reliance on the generalized quantum Stein’s lemma.
We further note that the additivity property of entanglement measures is more fundamental than asymptotic continuity for obtaining reversible manipulations of entangled states. Specifically, any additive entanglement measure allows for the derivation of a second law for zero-error transformations, i.e., transformations that convert $n$ copies of $\rho$ into $m$ copies of $\sigma$ exactly. Logarithmic negativity is an example of a measure that is additive and yet not asymptotically continuous [28, 29, 7]. We demonstrate that this measure leads to a theory with bounded entanglement distillation rates, even if an error margin is allowed in the asymptotic transformations. Additionally, we illustrate a phenomenon termed self-dilution: the asymptotic conversion of $n$ copies of an entangled state into $m>n$ copies of itself, with an error that can be made arbitrarily small in the asymptotic limit. More details with concrete examples are discussed in the Methods section.
@@
We will now discuss the difference of our methods from the approach considered in Ref. [7]. Specifically, Ref. [7] posits that reversible manipulations of entangled states are unattainable in frameworks reliant on non-entangling operations. In such frameworks, Alice and Bob, possessing many copies of a state $\rho$, are restricted to quantum operations that cannot generate entanglement from separable states. Under these conditions, the authors of [7] demonstrate the impossibility of reversible interconversion between certain entangled states $\rho$ and $\sigma$, i.e., $R(\rho\rightarrow\sigma)\times R(\sigma\rightarrow\rho)<1$. It is important to note that the methods in Ref. [7] do not cover the transformations considered in this work, because we allow the agents to further optimize on the operation in a way that depends on the system’s initial and final states. This is critical in enabling reversible interconversions within our framework, and moreover is a standard feature of catalytic processes [15, 16, 17]. Additionally, it is crucial to clarify that in our setup, Alice and Bob are also precluded from creating entanglement from separable states if the entanglement measure $E$ is additive and vanishes solely on separable states. These conditions are satisfied, for example, by the squashed entanglement [31, 27, 32].
@@
In the preceding discussion, we operated under the assumption that the main system and the battery return to an uncorrelated state at the procedure’s conclusion. However, this requirement can be relaxed to allow for correlations between them, with the constraint that the amount of entanglement in the battery does not decrease. In this scenario, any entanglement measure $E$ yields a reversible theory provided it satisfies four conditions: monotonicity under LOCC, continuity, additivity on product states, and general superadditivity, expressed as $E(\rho^{AA^{\prime}BB^{\prime}})\geq E(\rho^{AB})+E(\rho^{A^{\prime}B^{\prime}})$. This means that the amount of entanglement present in a joint state is always at least as high as the individual entanglement. Superadditivity has been shown to be a critical feature of resource monotones in general [33, 34, 17]. Notably, for pure bipartite states, transformations in this framework are entirely governed by the entanglement entropy, regardless of the entanglement measure chosen. This mirrors the uniqueness of entanglement entropy as the measure governing the transformations under standard LOCC [35, 8, 36]. For an in-depth discussion, we refer to the Supplemental Material.
@@
As elaborated upon in the Supplemental Material, not all entanglement measures prove useful within this context. Specifically, for certain measures such as geometric entanglement, the constraint in Eq. (2) fails to impose limitations on potential transformations, permitting arbitrary amplification of entanglement within the main system. This phenomenon bears resemblance to the concept of entanglement embezzlement, as previously discussed in [14].
Quantum thermodynamics
@@
The framework described in this work has immediate implications beyond entanglement theory. The second law of thermodynamics says that state transformations in classical thermodynamics are governed by the free energy [9]. However, this statement relies on several assumptions such as the negligibility of fluctuations and energetic coherence, which might not necessarily hold in the quantum regime. In particular, one commonly refers to quantum states without any coherence in the energy eigenbasis as energy-incoherent states.
@@
A commonly used model to study the thermodynamics of quantum systems is the framework of thermal operations, which explicitly model interactions with a thermal bath through an energy-preserving unitary [37]. For energy-incoherent states, the transformations are fully governed by a family of inequalities called the thermomajorization conditions [38]. The original form of the second law is subsequently recovered, if we consider catalytic transformations that allows correlations between the system and catalyst [39]. More precisely, it was shown that for any two energy-incoherent states $\rho,\sigma$, there exists a thermal operation that transforms $\rho^{S}$ into $\sigma^{S^{\prime}}$ if and only if $F(\rho^{S},H^{S})\geq F(\sigma^{S^{\prime}},H^{S^{\prime}})$, where $F$ is the free energy, defined as
@@
Here, $T$ is the temperature $Z$ is the partition function, and $\gamma=\exp\pqty{-H/k_{B}T}/Z$ is the thermal state associated with the system (that is characterized by $H$). Numerical evidences even suggest that low-dimensional catalysts already provide significance advantage [40], when it comes to the problem of simplifying the required unitary control over system and bath.
The obvious enhancement of catalysis in quantum thermodynamics even for energy-incoherent state transformations raises the following question: is it possible to extend this result to coherent states? In fact, it has recently been shown that the presence of catalysts in thermodynamics fully bridge between a hierarchy of thermal processes [41] on the level of the generated set of quantum channels, i.e. including energy-coherent state transformations. Nevertheless, it is still an open question whether such catalytic power also closes the gap between thermal operations and Gibbs-preserving ones. The most recent progress on this problem is Ref. [42], which suggests that there are Gibbs-preserving operations that require infinite coherence to be implemented; however, it is unclear whether this coherence can be used catalytically. One of the challenges to this open problem lies again in the challenge of retaining external correlations (or similarly, coherence) in the catalyst after the process, as exemplified in [43].
We answer this question by showing that the free energy determines general state transformations under thermal operations when we allow access to a thermodynamic battery. This framework includes catalysis as a special case, but relaxes the requirement of restoring the catalyst, and simply focuses on the reusability of the battery – requiring that the thermodynamic resource of the battery does not decrease. This approach is much more flexible and avoids the problems in catalysis such as fine-tuning and embezzling. More formally, we study transformations of the form $\rho^{S}\otimes\tau^{B}\to\sigma^{S^{\prime}}\otimes\tilde{\tau}^{B^{\prime}}$, along with the requirement that the free energy does not decrease $F(\tilde{\tau}^{B^{\prime}},H^{B^{\prime}})\geq F(\tau^{B},H^{B})$. Intuitively, this allows the battery to act as a source of coherence without giving away free energy. By analogous arguments to Theorem 1, we can show that at any finite temperature $T<\infty$, the allowed set of transformations is governed by free energy, i.e. $\rho^{S}\to\sigma^{S^{\prime}}$ if and only if $F(\rho^{S},H^{S})\geq F(\sigma^{S^{\prime}},H^{S^{\prime}})$.

### v2

@@ Abstract
A central question since the beginning of quantum information science is how two distant parties can convert one entangled state into another. It has been conjectured that such conversions could be executed reversibly in an asymptotic regime, mirroring the reversible nature of Carnot cycles in classical thermodynamics. While a conclusive proof of this conjecture has been missing so far, earlier studies have excluded reversible entanglement manipulation in various settings. In this work, we show that arbitrary mixed state entanglement transformations can be made reversible under local operations and classical communication, when assisted by an entanglement battery–an auxiliary quantum system that stores and supplies entanglement in a way that ensures no net entanglement is lost. In particular, the rate of transformation in the asymptotic limit can be quantitatively expressed as a ratio of entanglement present within the quantum states involved. Our setting allows to consider different entanglement quantifiers which give rise to unique principles governing state transformations, effectively constituting diverse manifestations of a “second law” of entanglement manipulation. These findings resolve a long-standing open question on the reversible manipulation of entangled states and are also applicable to multipartite entanglement and other quantum resource theories, including quantum thermodynamics.
[…]
The striking similarity between entanglement theory and thermodynamics naturally leads to the question: Does there exist a “second law of entanglement manipulation” [1], akin to its thermodynamic counterpart? In other words, is there an “entropy-like” quantity that governs all state transformations of entangled systems? This question is conceptually tied to the feasibility of reversibility, specifically in the sense of zero loss rate in the asymptotic limit. Let us denote the maximal rate of transformation of $n$ copies of $\rho$ to $m$ copies of $\sigma$ as $R(\rho\to\sigma)=\sup\Bqty{m/n}$, where the supremum is taken over a set of all allowed protocols of interest [9, Section 11.5]. When two states $\rho,\sigma$ can be transformed reversibly, then $R(\rho\to\sigma)R(\sigma\to\rho)=1$. Analogously to Carnot’s theorem, reversibility is connected to the efficiency of conversion between any two states. To see this, let us define the loss rate in a cycle $\rho\to\sigma\to\rho$ as $\ell(\rho\to\sigma\to\rho)=1-R(\rho\to\sigma)R(\sigma\to\rho)$. Then, a given protocol converts two states losslessly if and only if it is reversible.
It is known that for any two pure states $\psi,\phi$, we have $R(\psi\to\phi)=S(\psi_{A})/S(\phi_{A})$ [7], and in particular $R(\psi\to\phi)R(\phi\to\psi)=1$. In other words, the transformation is reversible. This implies that the distillation rate of singlets determines all transformation rates $R(\psi\to\phi)=R(\psi\to\psi^{-})/R(\phi\to\psi^{-})$, giving us a “second law of entanglement manipulation” for pure states. However, the situation no longer holds for mixed states. Extreme examples are the well-studied bound entangled states [10], where entanglement is required for their formation (under LOCC), yet one cannot extract any usable entanglement from these states in the form of singlets [10, 11].
To obtain a second law for entanglement beyond pure states, one has to move beyond the traditional LOCC paradigm. Over the past decade, a prominent approach was to connect the problem of state transformations to that of hypothesis testing [2]. In this framework, all protocols that inject a small amount of entanglement into the system were allowed, provided that this contribution vanishes in the asymptotic limit. Nonetheless, it is important to note that the assertions in [2] rely on a landmark result known as the generalized quantum Stein’s lemma [3]. Recent scrutiny [12] has raised questions about the validity of this lemma, casting uncertainty on the robustness of the conclusions drawn from its application.
Ref. [5] motivated recent renewed interest in the problem of (ir)reversibility. It was shown that when restricting to operations that preserve the set of separable states—known as non-entangling operations—then entanglement manipulation is necessarily irreversible. Even when we boost LOCC with catalysis, reversibility cannot be obtained due to the existence of bound entanglement in this setting [13, 14]. One can in principle recover reversibility by allowing the use of non-CPTP maps [15]. However, these maps are not physical and any implementation must involve classical postprocessing that comes with an associated sampling overhead. A different approach shows reversibility by adopting probabilistic protocols [16], which necessarily changes the meaning of reversibility. Indeed, in this we are only guaranteed that there is a non-zero chance that we recover the initial state. Are these sacrifices necessary to obtain a reversible theory of entanglement?
@@
In this article, we contribute to the above ongoing debate, by demonstrating the existence of a physically meaningful framework in which entanglement reversibility can be achieved. We introduce the concept of entanglement batteries [17] into our framework. An entanglement battery is an additional entangled system shared between Alice and Bob. In order to prevent the embezzling of resources [18], the battery must be returned at the end of the procedure with at least the same amount of entanglement. This idea can be viewed as a generalization of entanglement catalysis [19], a topic that has garnered significant interest recently [20, 21]. An entangled catalyst, in this context, is an ancillary quantum system in an entangled state provided to Alice and Bob. They are allowed to employ this catalyst in their transformation process, with the requirement that it must be returned in its original state. Our approach extends this setup by allowing changes in the state of the ancillary system, provided that there is no reduction in its entanglement.
@@
Reversible entanglement manipulations with entanglement battery. We demonstrate that reversible manipulations across all entangled states become feasible when an entanglement battery is incorporated. Specifically, we consider a setting where Alice and Bob can perform LOCC, and additionally are given access to a supplementary shared entangled state, which is our entanglement battery. This setting has been previously studied in Ref. [17], where it was shown that we can formulate a fluctuation theorem in pure state entanglement transformations with an explicit battery model. In this article, we generalize this framework to allow for a general battery and we study transformations between general mixed states. We show that we obtain a reversible theory when we impose the following restriction on the battery: Alice and Bob must not decrease the level of entanglement within the battery throughout the process. The setup is shown in Fig. 1.
@@
The basic property of any entanglement measure $E$ is that it should not increase under LOCC, i.e., $E(\Lambda[\rho])\leq E(\rho)$ for any LOCC protocol $\Lambda$ [22]. A particularly interesting class of entanglement measures for our analysis are quantifiers which are additive, i.e., they satisfy $E(\rho\otimes\tau)=E(\rho)+E(\tau)$ for any states $\rho$ and $\tau$. Additionally, we will consider entanglement quantifiers that exhibit asymptotic continuity [23]. This property implies that for two states close to each other in trace distance, the difference in the amount of entanglement can grow at most logarithmically in the Hilbert space dimension. We further require the entanglement measure to be finite for all states. The subsequent theorem offers a comprehensive characterization of state transformations within this setup.
@@
We provide a concise outline of our proof, with comprehensive details available in the End Matters. First, we leverage the properties of additive entanglement measures to establish that under the framework of LOCC augmented with an entanglement battery, the amount of entanglement does not increase. Subsequently, we introduce a protocol that transforms the state $\rho$ into $\sigma$ for any pair of states that comply with the stipulations outlined in Eq. (3). This protocol is remarkably straightforward: the battery is initialized in the target state $\sigma$, and then local transformations are employed to interchange the states of the main system and the battery. A similar technique has been used recently in [21] in the context of correlated catalysis, see also [17].
@@
An outline of the proof is provided here, with detailed elaboration available in the End Matters. Employing the properties of additive and asymptotically continuous entanglement measures, we first demonstrate that the transformation rate is upper-bounded by $E(\rho)/E(\sigma)$. The converse is shown by introducing a protocol that accomplishes the transformation at the aforementioned rate. Echoing the approach of Theorem 1, the optimal protocol involves a battery in an entangled state, specifically comprising copies of the target state $\sigma$. The transformation $\rho\rightarrow\sigma$ at the rate of $E(\rho)/E(\sigma)$ can be realized by permuting the main system with the battery.
An immediate consequence of our work is the construction of a framework of reversible entanglement manipulation. More specifically, we establish that $R(\rho\rightarrow\sigma)R(\sigma\rightarrow\rho)=1$ for any $\rho$ and $\sigma$, implying that in the asymptotic limit, any two entangled states $\rho$ and $\sigma$ can be interconverted reversibly. Furthermore, when the ratio $E(\rho)/E(\sigma)$ is rational, reversible interconversion is feasible even with a finite number of copies. We further notice that the optimal conversion rate can be achieved through a protocol involving local operations alone, which interestingly obviates the need for classical communication in this task. We suspect that this is an artifact of allowing the battery to contain any finite amount of entanglement. However, we anticipate that achieving the conversion with minimal entanglement in the battery may require some level of classical communication. In addition, we can argue that this framework is tight with respect to reversibility; in the Supplemental Material we show that if we require that the entanglement in the battery does not decrease when measured by two non-equivalent measures, then we cannot obtain a reversible framework.
@@
with the quantum conditional mutual information $I(A;B|E)$. We refer to the Supplemental Material for more details about the properties of squashed entanglement. We note that the squashed entanglement is not the only entanglement quantifier having these properties, another example is given in the Supplemental Material.
Our findings not only affirm the existence of a second law for entangled state transformations, but also link it to the squashed entanglement present in the involved quantum states. Prior to our work, it was hypothesized [2, 1] that should reversible manipulation of entangled states be achievable in any framework, the rates at which such reversible transformations occur would be linked to a different measure of entanglement: the regularized relative entropy of entanglement. This quantity is defined as [22]
@@
As detailed in the Supplementary Materials, we can extend our methodologies to establish a version of the second law of entanglement manipulations predicated on the relative entropy of entanglement. Additionally, we illustrate a phenomenon termed self-dilution: the asymptotic conversion of $n$ copies of an entangled state into $m>n$ copies of itself, with an error that can be made arbitrarily small in the asymptotic limit. This is possible when we use an additive but not asymptotically continuous entanglement measure such as logarithmic negativity. More details with concrete examples are discussed in the Supplemental Material.
@@
We will now discuss the difference of our methods from the approach considered in Ref. [5]. Specifically, Ref. [5] posits that reversible manipulations of entangled states are unattainable in frameworks reliant on non-entangling operations. In such frameworks, Alice and Bob, possessing many copies of a state $\rho$, are restricted to quantum operations that cannot generate entanglement from separable states. Under these conditions, the authors of [5] demonstrate the impossibility of reversible interconversion between certain entangled states $\rho$ and $\sigma$, i.e., $R(\rho\rightarrow\sigma)R(\sigma\rightarrow\rho)<1$. It is important to note that the methods in Ref. [5] do not cover the transformations considered in this work, because we allow the agents to have access to an additional system. This opens a degree of freedom since Alice and Bob can further optimize the protocol (i.e. operation and battery) in a way that depends on the system’s initial and final states. This is critical in enabling reversible interconversions within our framework, and moreover is a standard feature of catalytic processes [19, 20, 21]. Additionally, it is crucial to clarify that in our setup, Alice and Bob are also precluded from creating entanglement from separable states if the entanglement measure $E$ is additive and vanishes solely on separable states. These conditions are satisfied, for example, by the squashed entanglement [26, 24, 27]. In summary, we are focusing on non-entangling state transformations instead of non-entangling operations.
@@
In the preceding discussion, we operated under the assumption that the main system and the battery return to an uncorrelated state at the procedure’s conclusion. However, this requirement can be relaxed to allow for correlations between them, with the constraint that the amount of entanglement in the battery does not decrease. In this scenario, any entanglement measure $E$ yields a reversible theory provided it satisfies four conditions: monotonicity under LOCC, continuity, additivity on product states, and general superadditivity, expressed as $E(\rho^{AA^{\prime}BB^{\prime}})\geq E(\rho^{AB})+E(\rho^{A^{\prime}B^{\prime}})$. Superadditivity has been shown to be a critical feature of resource monotones in general [28, 29, 21]. Notably, for pure bipartite states, transformations in this framework are entirely governed by the entanglement entropy, regardless of the entanglement measure chosen. This mirrors the uniqueness of entanglement entropy as the measure governing the transformations under standard LOCC [30, 7, 31]. For an in-depth discussion, we refer to the Supplemental Material.
@@
Not all entanglement measures prove useful within this context. Specifically, for certain measures such as geometric entanglement, the constraint in Eq. (2) fails to impose limitations on potential transformations, permitting arbitrary amplification of entanglement within the main system. This phenomenon bears resemblance to the concept of entanglement embezzlement, previously discussed in [18]. We refer the interested reader to the Supplemental Material for more details.
@@
Quantum thermodynamics. The framework described in this work has immediate implications beyond entanglement theory. The second law of thermodynamics says that state transformations in classical thermodynamics are governed by the free energy [8]. However, this statement relies on several assumptions such as the negligibility of fluctuations and energetic coherence, which might not necessarily hold in the quantum regime. In particular, one commonly refers to quantum states without any coherence in the energy eigenbasis as energy-incoherent states.
@@
A commonly used model to study the thermodynamics of quantum systems is the framework of thermal operations, which explicitly models interactions with a thermal bath through an energy-preserving unitary [32]. For energy-incoherent states, the transformations are fully governed by a family of inequalities called the thermomajorization conditions [33]. The original form of the second law is subsequently recovered, if we consider catalytic transformations that allows correlations between the system and catalyst [34]. More precisely, it was shown that for any two energy-incoherent states $\rho,\sigma$, there exists a thermal operation that transforms $\rho^{S}$ into $\sigma^{S^{\prime}}$ if and only if $F(\rho^{S},H^{S})\geq F(\sigma^{S^{\prime}},H^{S^{\prime}})$, where $F$ is the free energy, defined as
@@
Here, $T$ is the temperature $Z$ is the partition function, and $\gamma=\exp\pqty{-H/k_{B}T}/Z$ is the thermal state associated with the system (that is characterized by $H$). Numerical evidences even suggest that low-dimensional catalysts already provide significant advantages [35], when it comes to the problem of simplifying the required unitary control over system and bath.
The obvious enhancement of catalysis in quantum thermodynamics even for energy-incoherent state transformations raises the following question: is it possible to extend the second law to coherent states? Previous results suggests that coherence and catalysis can interact in unexpected ways [36, 37, 38, 39, 40], and thus it is unclear if such an extension is possible. We answer this question by showing that the free energy determines general state transformations under thermal operations when we allow access to a thermodynamic battery. Formally, we study transformations of the form $\rho^{S}\otimes\tau^{B}\to\sigma^{S^{\prime}}\otimes\tilde{\tau}^{B^{\prime}}$, along with the requirement that the free energy does not decrease $F(\tilde{\tau}^{B^{\prime}},H^{B^{\prime}})\geq F(\tau^{B},H^{B})$. Intuitively, this allows the battery to act as a source of coherence without simultaneously giving away free energy. By analogous arguments to Theorem 1, we can show that at any finite temperature $T<\infty$, the allowed set of transformations is governed by free energy, i.e. $\rho^{S}\to\sigma^{S^{\prime}}$ if and only if $F(\rho^{S},H^{S})\geq F(\sigma^{S^{\prime}},H^{S^{\prime}})$.